OpenStudy (anonymous):

Solve. x^2 – 4x = 14

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

these are the answers given to choose from 2 ± radical 14 2 ± 2 radical 3 2 ± 3 radical 2 2 ± radical 10

OpenStudy (anonymous):

Possible intermediate steps: x^2-4 x = 14 Add 4 to both sides: x^2-4 x+4 = 18 Factor the left hand side: (x-2)^2 = 18 Take the square root of both sides: abs(x-2) = 3 sqrt(2) Eliminate the absolute value: x-2 = -3 sqrt(2) or x-2 = 3 sqrt(2) Add 2 to both sides: x = 2-3 sqrt(2) or x-2 = 3 sqrt(2) Add 2 to both sides: x = 2-3 sqrt(2) or x = 2+3 sqrt(2

OpenStudy (lalaly):

the second

OpenStudy (anonymous):

thankx guyz